OpenStudy (anonymous):

when a single card is drawn from a 52 card deck, find the odds in favor of getting a black 3 or a red 9

OpenStudy (anonymous):

I'd say 4/52 now simplify and get 1/13

OpenStudy (cwrw238):

1/52 + 1/52

OpenStudy (anonymous):

What cwrw238? There are 2 black 3's and 2 red 9's.

OpenStudy (anonymous):

there are two black 3's and two red 9's, so any one of those 4 cards will work. 4 good choices out of 52 possible choices = 4/52 = 1/13

OpenStudy (cwrw238):

yea - i stand corrected....

OpenStudy (anonymous):

got it. thanks!

OpenStudy (cwrw238):

i should play more card games!!!

OpenStudy (anonymous):

Haha, sorry! I thought you knew something I didn't about these types of problems!

OpenStudy (cwrw238):

thats ok